# Tomographic constraint on anisotropic cosmic birefringence

### v1

@@ IV Results
Fig. 4 shows the cosmic birefringence power spectrum measured from the Planck PR4 polarization data, with error bars derived from the standard simulations. In our baseline case, we employ the polarization map obtained from the Commander method for Galactic foreground cleaning and use the Planck $70\%$ Galactic mask. At large-angular scales, a nonzero isotropic birefringence can introduce a mean-field bias Namikawa:2020:biref . To mitigate the impact of the isotropic birefringence of $\sim 0.3\,$deg, we apply a uniform derotation of $\beta=-0.3$ degrees to the polarization map prior to reconstruction. At large angular scales, the Galactic foreground can also potentially affect the measurement. Therefore, we also measure the angular power spectrum in alternative scenarios, varying the Galactic mask, the isotropic derotation angle $\beta$, and the foreground cleaning method. In all cases, we compute the $\chi^{2}$-PTE of the measured power spectrum for multipoles in the range $1\leq L\leq 20$, finding that the measured spectrum is consistent with a null result.
@@
Here, $C_{L}^{\alpha\alpha,{\rm fid}}=2\pi/L(L+1)$, $C_{L}^{\rm f}$ corresponds to $C_{L}^{\hat{\alpha}\hat{\alpha},\rm th}$ evaluated at $A_{\rm CB}^{\rm rei}=0$, and ${\rm C}_{LL^{\prime}}$ is the covariance matrix of $C_{L}^{\rm f}$. We show $\mathcal{L}$ as a function of $A_{\rm CB}^{\rm rei}$ in Fig. 4. From this analysis, we find an upper limit of $A_{\rm CB}^{\rm rei}<1.3\times 10^{-2}$ at the $1\,\sigma$ level. It is important to note that the model proposed by Ref. Ferreira:2023 , which predicts $A_{\rm CB}^{\rm rei}\simeq 10^{-5}$, is still consistent with current observational constraints. However, this model would be excluded if we were to naively apply the anisotropic cosmic birefringence power spectrum measured in the previous studies.
@@ V Summary
Figure 5: Constraints on the amplitude of the scale-invariant power spectrum, $A_{\rm CB}^{\rm rei}$, derived from the cosmic birefringence power spectrum, shown as a function of polarization noise after Galactic foreground cleaning. We present results for both the delensed and non-delensed cases. The horizontal line represents the prediction from Ref. Ferreira:2023 , while the vertical lines indicate the expected noise levels for the PICO and LiteBIRD experiments.
We explored tomographic constraints on anisotropic cosmic birefringence using CMB polarization data. Specifically, we measured the angular power spectrum of anisotropic cosmic birefringence generated during the reionization epoch, using the Planck PR4 polarization maps. Our analysis found the power spectrum to be consistent with null results, leading to an upper bound on the amplitude of the scale-invariant power spectrum: $A_{\rm CB}^{\rm rei}<1.3\times 10^{-2}\,(1\,\sigma)$.
@@
where $C_{L}^{\alpha\alpha,\rm fid}=2\pi/L(L+1)$ and $f_{\rm sky}$ is a fraction of observed sky. The normalization given in Eq. (10), $A_{L}$, corresponds to the reconstruction noise power spectrum. The observed $E$- and $B$-mode power spectra are modeled as the sum of the lensed CMB and white-noise power spectra. We assume that, after Galactic foreground cleaning, the noise level is characterized by white noise with an amplitude $\sigma_{\rm P,post}$. Additionally, we account for the delensing of $B$-mode polarization, which reduces the cosmic variance from lensing and enhances sensitivity to primary $B$-modes Kesden:2002ku ; Seljak:2003pn . Specifically, we consider an $80\%$ delensing scenario for the space-based experiments, achievable through a combination of future ground-based CMB experiments and external mass tracers CMBS4 ; Namikawa:2023:LB-delens . In this scenario, the lensing $B$-mode power spectrum is scaled by a factor of $0.2$ before calculating $A_{L}$ in Eq. (10). LiteBIRD and PICO will improve the sensitivity to $A_{\rm CB}^{\rm rei}$ by more than three orders of magnitude compared to Planck, enabling them to test the model proposed by Ref. Ferreira:2023 .

### v2

@@ IV Results
Figure 4 shows the cosmic birefringence power spectrum measured from the Planck PR4 polarization data, with error bars derived from the standard simulations. In our baseline case, we employ the polarization map obtained from the Commander method for Galactic foreground cleaning and use the Planck $70\%$ Galactic mask. At large-angular scales, a nonzero isotropic birefringence can introduce a mean-field bias Namikawa:2020:biref . To mitigate the impact of a potential isotropic birefringence angle $\beta$, we apply a uniform derotation of $-\beta$ to the polarization map prior to reconstruction. We do not find any relevant difference in the measured birefringence spectrum for different values of $\beta=-0.1/0.3/0.5$ degrees. At large angular scales, the Galactic foreground can also potentially affect the measurement. Therefore, we also measure the angular power spectrum by varying the Galactic mask and the foreground cleaning method. In all cases, we compute the $\chi^{2}$-PTE of the measured power spectrum for multipoles in the range $1\leq L\leq 20$, finding that the measured spectrum is consistent with a null result.
@@
Here, $C_{L}^{\alpha\alpha,{\rm fid}}=2\pi/L(L+1)$, $C_{L}^{\rm f}$ corresponds to $C_{L}^{\hat{\alpha}\hat{\alpha},\rm th}$ evaluated at $A_{\rm CB}^{\rm rei}=0$, and ${\rm C}_{LL^{\prime}}$ is the covariance matrix of $C_{L}^{\rm f}$. We show $\mathcal{L}$ as a function of $A_{\rm CB}^{\rm rei}$ in Fig. 4. From this analysis, we find an upper limit of $A_{\rm CB}^{\rm rei}<1.3\times 10^{-2}$ at the $1\,\sigma$ level. It is important to note that the model proposed by Ref. Ferreira:2023 , which predicts $A_{\rm CB}^{\rm rei}\simeq 10^{-5}$, is still consistent with current observational constraints. However, this model would be excluded if we were to naively apply the anisotropic cosmic birefringence power spectrum measured in the previous studies Gluscevic:2012 ; PB15:rot ; BICEP2:2017lpa ; Contreras:2017 ; Namikawa:2020:biref ; SPT:2020:biref ; Gruppuso:2020 ; Bortolami:2022whx ; BK-LoS:2023 ; Zagatti:2024 .
@@ V Summary
Figure 5: Uncertainty on the amplitude of the scale-invariant power spectrum, $A_{\rm CB}^{\rm rei}$, derived from the cosmic birefringence power spectrum, shown as a function of polarization noise after Galactic foreground cleaning. We present results for both the delensed and non-delensed cases. The black horizontal line represents the prediction from Ref. Ferreira:2023 , while the vertical lines indicate the expected noise levels for the PICO and LiteBIRD experiments.
We explored tomographic constraints on anisotropic cosmic birefringence using CMB polarization data. Specifically, we measured the angular power spectrum of anisotropic cosmic birefringence generated during the reionization epoch, using the Planck PR4 polarization maps. Our analysis found the power spectrum to be consistent with null, leading to an upper bound on the amplitude of the scale-invariant power spectrum: $A_{\rm CB}^{\rm rei}<1.3\times 10^{-2}\,(1\,\sigma)$.
@@
where $C_{L}^{\alpha\alpha,\rm fid}=2\pi/L(L+1)$ and $f_{\rm sky}$ is the fraction of observed sky. The normalization given in Eq. (10), $A_{L}$, corresponds to the reconstruction noise power spectrum. The observed $E$- and $B$-mode power spectra are modeled as the sum of the lensed CMB and white-noise power spectra. We assume that, after Galactic foreground cleaning, the noise level is characterized by white noise with an amplitude $\sigma_{\rm P,post}$. Additionally, we account for the delensing of $B$-mode polarization, which reduces the cosmic variance from lensing and enhances sensitivity to primary $B$-modes Kesden:2002ku ; Seljak:2003pn . Specifically, we consider an $80\%$ delensing scenario for the space-based experiments, achievable through a combination of future ground-based CMB experiments and external mass tracers CMBS4 ; Namikawa:2023:LB-delens . In this scenario, the lensing $B$-mode power spectrum is scaled by a factor of $0.2$ before calculating $A_{L}$ in Eq. (10). LiteBIRD and PICO will improve the sensitivity to $A_{\rm CB}^{\rm rei}$ by more than three orders of magnitude compared to Planck, enabling them to test the model proposed by Ref. Ferreira:2023 .